OpenStudy (anonymous):

Find the 5 critical points for the following function: y=4sinpiex

OpenStudy (anonymous):

what is the function \[y=4\sin \pi e^x \]

OpenStudy (anonymous):

\[y=4\sin \pi ex\]

OpenStudy (anonymous):

without the "e" yes

OpenStudy (anonymous):

oh okay take the derivative \[y'=4 \pi \cos \pi x\]

OpenStudy (anonymous):

\[\cos (\pi x)=0\]